Direct Materials and Direct Labor Variances At the beginning of June, Kimber Toy Company budgeted 12,000 toy action figures tFlexible Overhead Budget Leno Manufacturing Company prepared the following factory overhead cost budget for the Press Departm16,000 Leno Manufacturing Company Factory Overhead Cost Budget-Press Department For the Month Ended November 30 Direct laborFactory Overhead Cost Variances The following data relate to factory overhead cost for the production of 5,000 computers: Actb. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favora
Answer #1
Direct materials and direct labor variances:
Direct material quantity variance=Standard rate of material*(Actual quantity consumed-Standard quantity)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Standard rate of material=$ 0.60 per pound
Standard quantity=Actual units produced*standard quantity required per pound
Standard quantity required per pound=Budgeted materials in pounds/Budgeted units to be produced
Budgeted materials in pounds=Budgeted direct materials/Standard rate of material=18000/0.60=30000 pounds
Standard quantity required per pound=30000/12000=2.5 pounds per unit
Standard quantity=10600*2.5=26500 pound
Actual quantity consumed=Actual direct material cost/Actual rate of material
There were no direct material price variance for June.
Hence, Standard rate of material=Actual rate of material=$0.60 per pound
Actual quantity consumed=16400/0.60=27333 pound
Direct material quantity variance=0.60*(27333-26500)=4999.8=500=$ 500 Unfavorable
Direct labor time variance=Standard rate of labor*(Actual hours worked-Standard hours)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Standard rate of labor=$14 per hour
Standard hours=Actual units produced*Standard hour per unit
Standard hours per unit=standard hours required/Budgeted units to be produced
Standard hours required=Budgeted direct labor cost/Standard rate of labor=11760/14=840 hours
Standard hours per unit=840/12000=0.07 hours
Standard hours=10600*0.07=742 hours
Actual hours worked=Actual direct labor cost/Actual rate of labor
There were no direct labor rate variance for June.
Hence, Standard rate of labor=Actual rate of labor=$14 per hour
Actual hours worked=10700/14=764 hours
Direct labor time variance=14*(764-840)=-1064=$ 1064 Favorable
Flexible overhead budget:
Variable overhead cost per unit=budgeted cost/Budgeted hours of production
Budgeted
cost
Budgeted
hours of
production
Variable
overhead
cost per
unit
a b a/b
Indirect labor 131600 14000 9.4
Power and light 5740 14000 0.41
Indirect materials 36400 14000 2.6
Fixed overhead cost will remain same at all levels of production
Factory overhead cost budget-Press department
Direct labor hours 12000 14000 16000
$ $ $
Variable overhead cost:
Indirect labor 112800 131600 150400
(12000*9.4) (14000*9.4) (16000*9.4)
Power and light 4920 5740 6560
(12000*0.41) (14000*0.41) (16000*0.41)
Indirect materials 31200 36400 41600
(12000*2.6) (14000*2.6) (16000*2.6)
Total variable factory overhead a 148920 173740 198560
Fixed factory overhead cost:
Supervisory salaries 60810 60810 60810
Depreciation of plant and equipment 38220 38220 38220
Insurance and property taxes 24320 24320 24320
Total fixed factory overhead b 123350 123350 123350
Total factory overhead cost a+b 272270 297090 321910
Factory overhead cost variances:
Variable factory overhead controllable variance=Actual variable factory overhead cost incurred-Budgeted variable factory overhead for 4000 hours
If the answer is negative, variance is favorable.Otherwise unfavorable.
Budgeted variable factory overhead for 5000 hours=(Budgeted hours)*(standard factory overhead rate per hour-Standard fixed factory overhead rate per hour)=5000*(26-5.25)=$ 103750
Variable factory overhead controllable variance=100600-103750=-3150=$ 3150 favorable
Fixed factory overhead volume variance=Productive capacity not used*Standard fixed factory overhead rate
If the answer is negative, variance is favorable.Otherwise unfavorable.
Productive capacity not used=Productive capacity at 100%-Budgeted level of production=8000-5000=3000 hours
Fixed factory overhead volume variance=3000*5.25=15750=$ 15750 Unfavorable
Total factory overhead cost variance=Variable factory overhead controllable variance+Fixed factory overhead volume variance
Total factory overhead cost variance=-3150+15750=12600=$ 12600 Unfavorable
Direct materials,Direct labor and factory overhead cost variance analysis:
a Direct materials price variance=Actual quantity of material used*(Actual rate of material-Standard rate of material)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Actual material used=7200 lbs.
Standard rate of material=$ 4.70 per pound
Actual rate of material=$ 4.50 per pound
Direct materials price variance=7200*(4.50-4.70)=-1440=$ 1440 (Favorable)
Direct Materials quantity variance=Standard rate of material*(Actual quantity consumed-Standard quantity)
If the answer is negative, variance is favorable.Otherwise unfavorable.
Standard rate of material=$ 4.70 per pound
Standard quantity=7300 lbs.
Actual material used=7200 lbs.
Material quantity variance=4.70*(7200-7300)=-470=$ 470 (Favorable)
Total direct materials cost variance=Direct materials price variance+Direct materials quantity variance
If the answer is negative, variance is favorable.Otherwise unfavorable.
Total material cost variance=-1440+-470=1910=$ 1910 (Favorable)